A US investor can invest $100 in the US or in Japan. The US interest rate is 4.5% and the Japanese interest rate is 3.1%. The current exchange rate is $1 US buys 12.5 Yen, and the investor believes the future exchange rate will be 12.2 Yen. Then the US investment yields a future value of ____ US dollars and the Japanese investment yields ____ US dollars.

a)102.5; 105.64

b)102.5; 103.50

c)104.5; 103.50

d)104.5; 105.64
Answer #1

Amount if invested in US = 100 * (1+0.045) = 100 * 1.045 = 104.50 USD

Spot rate 1 US = 12.5 JPY

100 USD = 12.5 * 100 = 1250 JPY

Amount after 1 year if invested in Japan = 1250 * (1+0.031) = 1250 * 1.031 = 1288.75 JPY

Converting JPY into USD = 1288.75 / 12.2 = 105.6352 ~ 105.64 USD

option d is the correct answer